OpenStudy (anonymous):

solve this equation. 2^(2x) + 2^(x+2) - 96 = 0 x=

OpenStudy (anonymous):

one can write 2^(2x) = (2^x)^2 say a=2^x then the equation becomes a^2+4a-96=0; a^2+12a-8a-96=0; a(a+12)-8(a+12)=0 (a-8)(a+12)=0 so a=8 or a=-12. solution #1. 2^x=8...2^x=2^3...x=3 solution #2. 2^x=-12. there is no solution for x in this case. SO the only solution is that x=3